# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ Abstract
This paper studies the relationship between the nuclear norm-regularized minimization problem, which minimizes the sum of a $C^{2}$ function $h$ and a positive multiple of the nuclear norm, and its factorized problem obtained by the Burer-Monteiro technique. We first prove that every second-order stationary point of the factorized problem corresponds to an approximate stationary point of its non-factorized counterpart, and those rank-deficient ones correspond to global minimizers of the latter problem when $h$ is additionally convex, conforming with the observations in [2, 15]. Next, discarding the rank condition on the second-order stationary points but assuming the convexity and Lipschitz differentiability of $h$, we characterize, with respect to some natural problem parameters, when every second-order stationary point of the factorized problem is a global minimizer of the corresponding nuclear norm-regularized problem. More precisely, we subdivide the class of Lipschitz differentiable convex $C^{2}$ functions into subclasses according to those natural parameters and characterize when each subclass consists solely of functions $h$ such that every second-order stationary point of the associated factorized model is a global minimizer of the nuclear norm regularized model. In particular, explicit counterexamples are established when the characterizing condition on the said parameters is violated.
@@ 1 Introduction
Low-rank matrix estimation has been an extremely important and versatile problem that has attracted intense research over the last two decades and found many applications across a wide range of domains, such as network science [12], machine learning [11, 26], quantum physics [22], control [25] and imaging [43, 13], to name a few. Natural formulations of the problem include the rank-constrained minimization problem (e.g., see [21, 46]):
$\begin{array}[]{rl}\displaystyle{\min_{X\in\mathbb{R}^{m\times n}}}&h(X)\\ {\rm s.t.}&\mathrm{rank}(X)\leq r,\end{array}$ (1.1)
the constrained rank minimization problem (e.g., see [8, 39]):
$\begin{array}[]{rl}\displaystyle{\min_{X\in\mathbb{R}^{m\times n}}}&\mathrm{rank}(X)\\ {\rm s.t.}&h(X)\leq c,\end{array}$ (1.2)
or the rank-regularized minimization problem (e.g., see [23, 20]):
$\min_{X\in\mathbb{R}^{m\times n}}h(X)+\lambda\cdot\mathrm{rank}(X).$ (1.3)
In the above formulations, $r,c,\lambda>0$ are constants prescribed by the modelers and $h$ is a function representing the misfit between the output predicted by $X$ and the true observations.
Unfortunately, due to the non-convexity and combinatorial nature of the rank function, the optimization problems (1.1), (1.2) and (1.3) are difficult to solve in general. For computational tractability, many convex and non-convex surrogates were proposed [33, 35, 27, 39]. This paper focuses on the following surrogate:
@@
where $h$ is assumed to be twice continuously differentiable for the theoretical analysis, $\lambda>0$, and $\|\cdot\|_{*}$ denotes the nuclear norm. Problem (1.4) can be seen as an approximation to problem (1.3). Indeed, it was shown that the nuclear norm is the convex envelope of the rank function [8]. The upshot of problem (1.4) is that it is in the form of the so-called composite minimization that has been heavily studied in the literature, especially when $h$ is convex. Therefore, in principle it can be solved by many existing algorithms for composite minimization, including in particular various proximal algorithms [31, 18, 38, 40] in view of the closed-form expression of the proximal operator of the nuclear norm [7].
Nonetheless, in contemporary applications, the dimensions $m$ and $n$ of the decision variable $X$ can potentially be extremely high, rendering existing methods inapplicable. For example, in collaborative filtering, which is a classical application of low-rank matrix estimation, the dimensions $m$ and $n$ could be of the order of millions or even higher [26]. Worse still, the computational cost of the proximal operator associated with the nuclear norm, which is a fundamental building block of many existing algorithms for solving problem (1.4), is a cubic function in $m$ and $n$, as it involves the singular value decomposition of $X$. To circumvent this, researchers proposed to solve problem (1.4) via the Burer-Monteiro factorization technique [5, 6, 32, 46, 44], which replaces the variable $X$ by a low-rank approximation $UV^{\top}$ and solves the resulting problem:
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm and $r$ is an integer parameter specified by the modeler. Any optimal solution $(U^{*},V^{*})$ of problem (1.5) corresponds to an approximately optimal solution $U^{*}{V^{*}}^{\top}$ to problem (1.4). The advantage of the factorized problem (1.5) over problem (1.4) is twofold. First, the objective function in the factorized problem (1.5) is differentiable as soon as $h$ is. In contrast, the objective function in problem (1.4) is nonsmooth because of the nuclear norm. Second, we often choose $r\ll\min\{m,n\}$ in practice. The total size $r(m+n)$ of the matrix variables $(U,V)$ is therefore substantially smaller than the size $mn$ of the variable $X$ in the non-factorized counterpart (1.4).
Since our goal is to solve problem (1.4), the factorization rank $r$ cannot be too small. Indeed, optimal solutions of problem (1.4) cannot be recovered by solving problem (1.5) through the correspondence $(U,V)\mapsto UV^{\top}$ if $r$ is less than the minimum rank $r^{*}$ of the optimal solutions of problem (1.4). This issue is currently addressed indirectly as follows. First, it can be readily shown that $U^{*}{V^{*}}^{\top}$ is a global minimizer of problem (1.4) for any global minimizer $(U^{*},V^{*})$ of problem (1.5) if $r\geq r^{*}$ (e.g., see [16, Lemma 1]). Second, despite the non-convexity due to the bilinear term $UV^{\top}$, the factorized problem (1.5) has no spurious local minimizers or second-order stationary points if $r$ is sufficiently large [19, 44], which implies that one can actually solve problem (1.4) with a convex $h$ by using any optimization algorithm with a second-order convergence guarantee on problem (1.5). However, a proper choice of the parameter $r$ is highly nontrivial. In particular, as demonstrated by an example constructed in [44], merely having $r\geq r^{*}$ is not enough in general, let alone that the minimum solution rank $r^{*}$ is often unknown in practice. Our paper also revolves around the choice of the factorization rank $r$ by asking a different but more direct question:
@@
With this definition, our problem is equivalent to the investigation of the $r$-factorizability of the objective function $f$ of problem (1.4).
Most of the existing works on this question focus on the unregularized counterpart:
@@
and rely on the restricted isometry property of $h$. Here, we recall that for $\delta>0$ and integers $s,t\geq 0$, a twice continuously differentiable function $h:\mathbb{R}^{m\times n}\to\mathbb{R}$ is said to satisfy $\delta$-$\mathrm{RIP}_{s,t}$ condition [19, 45, 42] if for all $X,H\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(X)\leq s$ and $\mathrm{rank}(H)\leq t$, it holds that
The state-of-the-art results were established in [42, Corollary 2], which showed that all second-order stationary points of problem (1.6) are global minimizers of problem (1.6) if $h$ satisfies the $\tfrac{1}{3}$-$\mathrm{RIP}_{2r,2r}$ condition and that there exists a function $h$ satisfying the $\tfrac{1}{2}$-$\mathrm{RIP}_{2r,2r}$ condition but possessing a spurious second-order stationary point. Other theoretical results can be found in [10, 9, 46]. Interestingly, the question about the choice of the factorization rank $r$ has also been investigated in the symmetric case, where the non-factorized and factorized problems are to minimize $h(X)$ and $h(UU^{\top})$, respectively, where $X$ is a symmetric positive semidefinite matrix. In the special case of semidefinite programming, it has been resolved rather satisfactorily [3, 4, 29].
@@
For asymmetric, regularized problem 1.4, fewer works have been done. A common assumption in these works is that there exists an optimal solution to problem (1.4) and $r$ is chosen to be at least the minimum solution rank $r^{*}$. In [24, Theorem 3], the author showed that if $h$ is convex quadratic and satisfies the $\delta$-$\mathrm{RIP}_{2r,2r}$ condition with $\delta<\tfrac{1}{3}$, then the corresponding $f$ in problem (1.4) is $r$-factorizable. A similar result was established in [19, Theorem 2] for a general twice continuously differentiable convex function $h$ with a more restrictive bound on $\delta$. Later, in [15, Theorem 1], [^1] The results in [15] were stated in terms of restricted strong convexity and restricted smoothness. The moduli $\alpha$ and $\beta$ therein correspond to $1-\delta$ and $1+\delta$ in our discussion here, respectively. it was shown that when $h$ satisfies the $\delta$-$\mathrm{RIP}_{2r,2r}$ condition with $\delta<\tfrac{1}{3}$, then the corresponding $f$ in problem (1.4) is $r$-factorizable, and when $\delta\geq\frac{1}{3}$, a second-order stationary point of problem (1.5) corresponds to an approximate stationary point of problem (1.4). Notice that all these results rely on RIP-type conditions. It remains unclear whether these conditions are necessary for the $r$-factorizability. It is also worth mentioning that verifying RIP-type conditions is NP-hard [37].
Unlike most of the existing works, we do not invoke any RIP-type assumptions. The work closest to our non-RIP approach is [44], which considered the following pair of optimization problems:
where $\mathbb{S}_{+}^{n}$ is the set of $n\times n$ symmetric matrices and $h$ is $C^{2}$, Lipschitz differentiable and strongly convex, and showed that all second-order stationary points $U^{*}$ of $\widetilde{h}$ satisfy that $U^{*}{U^{*}}^{\top}$ is the unique minimizer $X^{*}$ of $h$ over ${\mathbb{S}}_{+}^{n}$ under suitable conditions on the factorization rank $r$, solution rank $\mathrm{rank}(X^{*})$ and the condition number $\kappa$ (i.e., the ratio between the Lipschitz constant of $\nabla h$ and the strong convexity modulus of $h$), namely (1) $r\geq{\rm rank}(X^{*})$ and $\kappa<3$; or (2) $n>r\geq{\rm rank}(X^{*})$ and $r>\tfrac{1}{4}(\kappa-1)^{2}{\rm rank}(X^{*})$; the author also constructed a function $h$ with $\kappa=3$ such that $\widetilde{h}$ has a second-order stationary point that does not correspond to any global minimizer of $h$ over $\mathbb{S}_{+}^{n}$. Our work can be seen as an extension of the studies in [44] to the asymmetric, regularized case, and is more general in the sense that we consider not only strongly but also non-strongly convex $h$. More precisely, instead of the problem pairs (1.7), we focus on the problems (1.4) and (1.5) and study the $r$-factorizability of $f$ in connection with a set of natural problem parameters, including the condition number $\kappa$ of $h$, the Lipschitz constant of $\nabla h$, the solution rank of problem (1.4) and the rank of a second-order stationary point of 1.5.
We now summarize our technical contributions. First, we prove that every second-order stationary point of problem (1.5) corresponds to an approximate stationary point of problem (1.4) via the mapping $(U,V)\mapsto UV^{\top}$. In particular, when the second-order stationary point is rank-deficient, it corresponds to a stationary point of problem (1.4), which is in accordance with the findings of [2, 15]. Consequently, if all second-order stationary points of problem (1.5) are rank-deficient and $h$ is convex, then $f$ is $r$-factorizable. Next, discarding the rank condition on the second-order stationary points but assuming the convexity and Lipschitz differentiability of $h$, we characterize, with respect to some natural problem parameters, when every second-order stationary point of the factorized problem (1.5) is a global minimizer of non-factorized problem (1.4). More precisely, we subdivide the class of Lipschitz differentiable convex $C^{2}$ functions into subclasses according to those natural parameters and characterize when each subclass consists solely of functions $h$ such that every second-order stationary point of problem (1.5) is a global minimizer of problem (1.4). Furthermore, explicit counterexamples are established when the characterizing condition on the said parameters is violated. To our knowledge, our results are the first characterizations of $r$-factorizability in terms of these natural parameters.

### v2

@@ Abstract
This paper studies the relationship between the nuclear norm-regularized minimization problem, which minimizes the sum of a $C^{2}$ function $h$ and a positive multiple of the nuclear norm, denoted by $f$, and its factorized problem obtained by the Burer-Monteiro technique. We are interested in deriving conditions that ensure every second-order stationary point of the factorized problem corresponds to a global minimizer of $f$, a property we call the $r$-factorizability of $f$ in this paper. Under suitable restricted isometry property (RIP) type assumptions on $h$, we prove the $r$-factorizability of $f$. Moreover, the RIP constant in our paper is tight, in the sense that we can construct concrete examples of $f$ that fail to be $r$-factorizable when the RIP constant is below the threshold. Our technique for constructing such examples is novel and may be of independent interest: specifically, we use a variant of the Von Neumann’s trace inequality and relate the existence of such examples to the optimal value of a quadratic program involving the RIP constant, then we explicitly solve this optimization problem to detect all the possible counterexamples.
@@ 1 Introduction
Low-rank matrix estimation has been an extremely important and versatile problem that has attracted intense research over the last two decades and found many applications across a wide range of domains, such as network science [13], machine learning [12, 24], quantum physics [21], control [23] and imaging [34, 14], to name but a few. This paper focuses on the following low-rank optimization problem:
@@
where $h$ is assumed to be twice continuously differentiable, $\lambda>0$, and $\|\cdot\|_{*}$ denotes the nuclear norm. Without loss of generality, we assume $m\leq n$ throughout the paper. Problem (1.1) takes the form of the so-called composite minimization which has been heavily studied, especially when $h$ is convex. Therefore, in principle it can be solved by many existing algorithms for composite minimization, including in particular various proximal algorithms [26, 18, 30, 31], thanks to the closed-form expression of the proximal operator of the nuclear norm [8].
Nonetheless, in contemporary applications, the dimensions $m$ and $n$ of the decision variable $X$ can potentially be extremely high, rendering methods working directly with the variable $X$ impractical. For example, in collaborative filtering, which is a classical application of low-rank matrix estimation, the dimensions $m$ and $n$ could be of the order of millions or even higher [24]. Worse still, the computational cost of the proximal operator associated with the nuclear norm, which is a fundamental building block of many existing algorithms for solving problem (1.1), is a cubic function in $m$ and $n$, as it involves the singular value decomposition of $X$. To circumvent this, researchers proposed to tackle problem (1.1) via the Burer-Monteiro factorization technique [6, 7, 27, 38, 35], which replaces the variable $X$ by a low-rank approximation $UV^{\top}$ and solves the resulting problem:
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm and $r\in[1,m]$ is an integer parameter specified by the modeler. The advantage of the factorized problem (1.2) over problem (1.1) is twofold. First, the objective function in the factorized problem (1.2) is differentiable as long as $h$ is. In contrast, the objective function in problem (1.1) is nonsmooth because of the nuclear norm. Second, we often choose $r\ll\min\{m,n\}$ in practice. The total size $r(m+n)$ of the matrix variables $(U,V)$ is therefore substantially smaller than the size $mn$ of the variable $X$ in the non-factorized counterpart (1.1).
Since our goal is to solve problem (1.1), the factorization rank $r$ in (1.2) cannot be too small. Indeed, optimal solutions of problem (1.1) cannot be recovered by solving problem (1.2) through the correspondence $(U,V)\mapsto UV^{\top}$ if $r$ is less than the minimum rank $r^{*}$ of the optimal solutions of problem (1.1). This issue is currently addressed indirectly as follows. First, it can be readily shown that $U^{*}{V^{*}}^{\top}$ is a global minimizer of problem (1.1) for any global minimizer $(U^{*},V^{*})$ of problem (1.2) if $r\geq r^{*}$ (e.g., see [17, Lemma 1]). Second, despite the non-convexity due to the bilinear term $UV^{\top}$, the factorized problem (1.2) has no spurious local minimizers or second-order stationary points if $r$ is sufficiently large and $h$ satisfies certain technical conditions [19, 35], which implies that one can actually solve problem (1.1) with a strongly convex $h$ by using any optimization algorithm with a second-order convergence guarantee on problem (1.2). However, a proper choice of the parameter $r$ is highly nontrivial. As demonstrated by an example constructed in [35], merely having $r\geq r^{*}$ is not enough in general, let alone that the minimum solution rank $r^{*}$ is often unknown in practice. Our paper also revolves around the choice of the factorization rank $r$ by asking a different but more direct question:
@@
With this definition, our problem is equivalent to the investigation of the $r$-factorizability of the objective function $f$ of problem (1.1). Most previous works studying the $r$-factorizability, if not all, rely on the restricted isometry property of $h$ [10, 9, 38, 33]. Here, we recall that for $\delta>0$ and integers $s,t\geq 0$, a twice continuously differentiable function $h:\mathbb{R}^{m\times n}\to\mathbb{R}$ is said to satisfy $\delta$-$\mathrm{RIP}_{s,t}$ condition [19, 37, 33] if for all $X,H\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(X)\leq s$ and $\mathrm{rank}(H)\leq t$, it holds that
A closely related subject concerns the factorizability in the symmetric, unregularized setting, in which case the analogue problem pairs are
where $\mathbb{S}_{+}^{n}$ is the set of $n\times n$ symmetric positive semidefinite matrices and $h$ is $C^{2}$. In [35, 36], the author considered the case where $\nabla h$ is Lipschitz differentiable and $h$ is strongly convex, and showed that all second-order stationary points $U^{*}$ of $\widetilde{h}$ satisfy that $U^{*}{U^{*}}^{\top}$ is the unique minimizer $X^{*}$ of $h$ over ${\mathbb{S}}_{+}^{n}$ under suitable conditions on the factorization rank $r$, solution rank $r^{*}=\mathrm{rank}(X^{*})$ and the condition number $\kappa$ (i.e., the ratio between the Lipschitz constant of $\nabla h$ and the strong convexity modulus of $h$), namely (1) $r\geq r^{*}$ and $\kappa<3$; or (2) $n>r\geq r^{*}$ and $r>\tfrac{1}{4}(\kappa-1)^{2}r^{*}$; the author also constructed a function $h$ with $\kappa=3$ such that $\widetilde{h}$ has a second-order stationary point that does not correspond to any global minimizer of $h$ over $\mathbb{S}_{+}^{n}$. The result was then extended to the class of non-strongly convex quadratic functions satisfying the $\delta$-$\mathrm{RIP}_{r+r^{*},r+r^{*}}$ condition in [35, Corollary 1.5] or [36, Theorem 1.4], with essentially the same bound on $\delta$.
The asymmetric, unregularized case has also been studied in the literature, where the factorized problem is
@@
It is known that $(U,V)\mapsto h(UV^{\top})$ may have spurious second-order stationary points as long as $r<m$ even when $h$ satisfies the (strongest) 0-$\mathrm{RIP}_{m,m}$ condition, as demonstrated by [36, Example 1.8]. This motivates the search of formulations that are equivalent to 1.4 but with better landscape properties. One example is
$\min_{U\in\mathbb{R}^{m\times r},V\in\mathbb{R}^{n\times r}}\tilde{F}(U,V):=h(UV^{\top})+\beta\|U^{\top}U-V^{\top}V\|_{F}^{2},$ (1.5)
which was first proposed in [28]. Upon defining $h_{a}:\mathbb{S}^{m+n}\to\mathbb{S}^{m+n}$ as
$h_{a}(X):=h(X_{2})+\beta(\|X_{1}\|_{F}^{2}+\|X_{3}\|_{F}^{2}-2\|X_{2}\|_{F}^{2}),~~X=\begin{bmatrix}X_{1}&X_{2}\\ X_{2}^{\top}&X_{3}\end{bmatrix},~X_{1}\in\mathbb{S}^{m},$ (1.6)
[…]
@@
For the assymetric, regularized problem 1.1, fewer works have been done. A common assumption in these works is that there exists an optimal solution to problem (1.1) and $r$ is chosen to be at least the minimum solution rank $r^{*}$. In [22, Theorem 3], the author showed that if $h$ is convex quadratic and satisfies the $\delta$-$\mathrm{RIP}_{2r,2r}$ condition with $\delta<\tfrac{1}{3}$, then the corresponding $f$ in problem (1.1) is $r$-factorizable. A similar result was established in [19, Theorem 2] for a general twice continuously differentiable convex function $h$ with a more restrictive bound on $\delta$. Later, in [16, Theorem 1], [^1] The results in [16] were stated in terms of restricted strong convexity and restricted smoothness. The moduli $\alpha$ and $\beta$ therein correspond to $1-\delta$ and $1+\delta$ in our discussion here, respectively. it was shown that when $h$ satisfies the $\delta$-$\mathrm{RIP}_{2r,2r}$ condition with $\delta<\tfrac{1}{3}$, then the corresponding $f$ in problem (1.1) is $r$-factorizable, and when $\delta\geq\frac{1}{3}$, a second-order stationary point of problem (1.2) corresponds to an approximate stationary point of problem (1.1). We are not aware of any tight result for 1.1, in the sense that once the proposed sufficient conditions fail, then an $h$ can be explicitly constructed with the corresponding $f$ in (1.1) being non-$r$-factorizable.
It is tempting to reduce the asymmetric problem to a symmetric problem, as in the unregularized case. Unfortunately, as we shall discuss in Section 1.1, this idea is inapplicable in the regularized case. To circumvent this, we adopt a different strategy to tackle the factorizability problem, which is outlined in Section 1.1. Our new techniques enable us to derive tight RIP-type conditions in the sense of 1.3 below.
To present our main results, we define the following classes.
Definition 1.2.
Let $L\in(0,\infty)$, $\mu>0$, $q,r^{*}\in[m]\cup\{0\}$. We define $\mathfrak{S}(L,\mu,q,r^{*})$ to be the set of all $h\in C^{2}(\mathbb{R}^{m\times n})$ satisfying the following conditions:
  1. (i)
    For all $X,H\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(X),\mathrm{rank}(H)\leq q+r^{*}$, it holds that:
    $\mu\|H\|_{F}^{2}\leq\nabla^{2}h(X)[H,H]\leq L\|H\|_{F}^{2}.$ (1.7)
  2. (ii)
    There exists a global minimizer $X^{*}\in\mathbb{R}^{m\times n}$ of $f$ in (1.1) satisfying $\mathrm{rank}(X^{*})=r^{*}$.
When $L=1+\delta$ and $\mu=1-\delta$, 1.7 can be viewed as the $\delta$-$\mathrm{RIP}_{q+r^{*},q+r^{*}}$ condition. Let us also note that when $q+r^{*}\geq m$, 1.7 reduces to the $\mu$-strong convexity of $h$ and $L$-Lipschitz continuity of $\nabla h$. To simplify the notation, we denote this latter kind of function classes by $\mathfrak{S}(L,\mu,r^{*})$. Our first main result is the following theorem, which characterizes the $r$-factorizability of $\mathfrak{S}(L,\mu,r^{*})$ in the following sense: when the conditions in 1.3 are satisfied, then for all $h$ in $\mathfrak{S}(L,\mu,r^{*})$ the corresponding $f$ in 1.1 is $r$-factorizable, and if not, there exists $h\in\mathfrak{S}(L,\mu,r^{*})$ such that the corresponding $f$ in 1.1 is not $r$-factorizable. [^2] We will refer to this property as “tight in the sense of 1.3” for the rest of this paper.
Theorem 1.3.
Let $r^{*}\in[m]\cup\{0\}$, $r\in[m]$, $\infty>L\geq\mu>0$ and $\kappa:=\frac{L}{\mu}\geq 1$. Suppose that $^{*}, rr$ and $\kappa$ satisfy any of the following conditions:
  1. (1)
    $r=m$.
  2. (2)
    $r\geq r^{*}$ and $\min\{r,m-r^{*}\}>\frac{(\kappa-1)^{2}}{4}\min\{r^{*},m-r\}$.
Then for all $h\in\mathfrak{S}(L,\mu,r^{*})$, the corresponding $f$ in 1.1 is $r$-factorizable. Otherwise, there exists a quadratic $h\in\mathfrak{S}(L,\mu,r^{*})$ such that the corresponding $f$ in 1.1 is not $r$-factorizable.
Next, by analyzing suitable subspaces, we also obtain the following corollary from Theorem 1.3.
Corollary 1.4.
1.3 holds when $\mathfrak{S}(L,\mu,r^{*})$ is replaced by $\mathfrak{S}(L,\mu,r,r^{*})$.
In the case of $r+r^{*}<m$, $L=1+\delta$ and $\mu=1-\delta$, Corollary 1.4 implies that $f$ is $r$-factorizable if $r/r^{*}>\frac{\delta^{2}}{(1-\delta)^{2}}$ and $h$ satisfies the $\delta$-$\mathrm{RIP}_{r+r^{*},r+r^{*}}$ condition with a global minimizer $X^{*}\in\mathbb{R}^{m\times n}$ satisfying $\mathrm{rank}(X^{*})=r^{*}$, and that an explicit counterexample can be constructed if $r/r^{*}\leq\frac{\delta^{2}}{(1-\delta)^{2}}$. This bound matches the necessity condition in [36, Theorem 1.6], which studied the different model 1.5. However, we would like to point out that the sufficient condition in [36, Theorem 1.6] requires $r/r^{*}>4\delta^{2}/(1-2\delta)^{2}$, which does not match the necessary condition there.
Interestingly, our bound here also matches the bound in [35, Theorem 1.1], which is for 1.3 though. However, as we will discuss in Section 1.1 below, our results do not follow from [35] and cannot imply the results in [35] and [36], since we require the constant $\lambda$ to be positive in 1.1, which yields special structures on the first-order stationary points of $F_{r}$ in 1.2, c.f. Proposition 3.2.
1.1 Sketch of proof techniques
To the best of our knowledge, all the previous works [28, 9, 33, 36] on the non-existence of spurious second-order stationary point of asymmetric problem 1.5 rely on symmetric reduction. To study the factorizability of 1.1, naturally, one would want to invoke a similar reduction and then invoke results for the symmetric case, such as [35, 36]. However, such an idea will not work, as we now explain. First, note that the symmetric case counterpart of 1.1 is to minimize
$h_{s}(X)=h(X_{2})+\frac{\lambda}{2}\mathrm{tr}(X),~~X=\begin{bmatrix}X_{1}&X_{2}\\ X_{2}^{\top}&X_{3}\end{bmatrix}$.
We therefore have $h_{s}\left(\begin{bmatrix}U\\ V\end{bmatrix}\begin{bmatrix}U^{\top}\!&\!V^{\top}\end{bmatrix}\!\right)=F_{r}(U,V)$. However, $h_{s}$ loses the $\mathrm{RIP}$ property by Fact 5.1, since $h_{s}$ is not even strongly convex on the one-dimensional linear subspace $\{tE_{11}:~t\in\mathbb{R}\}$, where $E_{11}\in\mathbb{R}^{m\times n}$ is the matrix whose $(1,1)$-entry is $1$ and is zero otherwise. Therefore, the classical techniques in the literature [28, 9, 33, 36] are not applicable here.
In contrast, our proofs do not rely on the symmetric reduction technique. We start with the following well-known fact, essentially a restatement of [25, Theorem 2.1.12]: if $h\in C^{1}(\mathbb{R}^{m\times n})$ is $\mu$-strongly convex and $\nabla h$ is $L$-Lipschitz continuous, then for all $X,Y\in\mathbb{R}^{m\times n}$ it holds that
$(L-\mu)\langle\nabla h(X)-\mu X-(\nabla h(Y)-\mu Y),X-Y\rangle\geq\|\nabla h(X)-\mu X-(\nabla h(Y)-\mu Y)\|_{F}^{2}.$ (1.8)
Next, by analyzing the first-order stationary point of $F_{r}$ in 1.2, we show that if $(\bar{U},\bar{V})$ is a stationary point of $F_{r}$, then $\bar{X}=\bar{U}\bar{V}^{\top}$ is a pseudo stationary point of $f$, in which case $\bar{X}$ and $\nabla h(\bar{X})$ admit simultaneous singular value decompositions; see Proposition 3.2 and Definition 3.4. Using a variant of Von Neumann’s trace inequality (see Lemma 2.3), we transform 1.8 into a bound concerning the singular values of $\bar{X},X^{*},\nabla h(\bar{X}),\nabla h(X^{*})$, when $X^{*}$ and $\bar{X}$ are pseudo stationary points of $f$. Then, the existence of counterexamples can be transformed to the existence of a certain feasible solution to a quadratic program involving $L$ and $\mu$; see 4.16. We then solve this optimization problem analytically, which yields tight bounds in the sense of 1.3.
The generalization to the RIP case is based on the observation that, when 1.7 holds, the function $h$ is $\mu$-strongly convex and $\nabla h$ is $L$-Lipschitz continuous on any linear subspace of $\mathbb{R}^{m\times n}$ consisting of merely matrices whose rank is no more than $q+r^{*}$; see Fact 5.1. In particular, for $\mathrm{rank}(X^{*})=r^{*}$ and $\mathrm{rank}(\bar{X})=r$, we can restrict the function on a linear subspace of $\mathbb{R}^{(r+r^{*})\times(r+r^{*})}$ under proper orthogonal transformation, which contains $X^{*}$ and $\bar{X}$ (see Lemma 5.2), and then apply the result for the strongly convex case (see Proposition 5.3).
Let us now compare our proof techniques to those used in the literature. In [35, 36], tight bounds on the $\mathrm{RIP}$ constant $\delta$ analogous to those in our 1.3 were obtained for 1.3, and their bounds are the same as ours despite the difference in the problems under study. Then the result in [36] for 1.3 was generalized to 1.5 by using symmetric reduction. Note that the best bounds on the $\mathrm{RIP}$ constant $\delta$ in [36, Theorem 1.6] for 1.5 are also not tight as in our Theorem 1.3, which may be attributed to their use of the reduction to the symmetric case, so that the $\delta$-$\mathrm{RIP}_{k,k}$ condition of $h$ only yields the $2\delta$-$\mathrm{RIP}_{k,k}$ condition on $h_{a}$ in 1.6. Although the tight bound in [35, 36] for 1.3 aligns with our bound, their techniques are inapplicable in our context. In [35], the existence of counterexamples was formulated as an semidefinite programming (SDP) problem, then the author solved the optimization problem analytically to get a tight RIP constant bound. In [36], the author showed that the existence of (quadratic) counterexamples can be equivalently formulated as the existence of escape direction, and a sharp escape direction is constructed explicitly based on the tangent-normal decomposition of the manifold of $\mathrm{rank}$-$r$ positive semidefinite matrices. In contrast, we formulate a quadratic programming problem and solve it analytically. Our proof techniques are also different from those in [19, 16, 22], which focus on 1.2. Indeed, our proof is based on analyzing the singular values of the global minimizer $X^{*}$ appeared in Definition 1.2 and other related matrices (see Proposition 4.4), which leads to the tight bounds in the sense of 1.3.